# ViClaim: A Multilingual Multilabel Dataset for Automatic Claim Detection in Videos

### v1

@@ 1 Introduction
In this work, we present the first step towards misinformation detection in videos by introducing ViClaim, a novel dataset where transcripts of YouTube Short videos are annotated using a custom annotation tool on a sentence level with the claim taxonomy introduced in Panchendrarajan and Zubiaga (2024). Each sentence is annotated to indicate whether it contains a claim, an opinion, or both, and if a claim is present, whether it is check-worthy. Since sentences may fall into multiple categories simultaneously, the task is framed as a multi-label classification problem. The dataset encompasses 1,798 videos in three languages (English, German, and Spanish) across six topics—five political and one entertainment—allowing for investigating domain transfer capabilities. Each sentence was annotated by four independent annotators, resulting in a total of 17,116 annotated sentences. Thus, ViClaim offers a rich database for future research on multimodal misinformation detection. As a first step to showcase the utility of ViClaim, we trained several baseline models on the transcripts (leaving multimodal models for future work). Our best-performing model achieved an F1 score of 89.9 for checkworthiness classification, 77.6 for detecting non-check-worthy claims, and 83.6 for opinion detection. These results demonstrate both the effectiveness of the dataset and the challenges posed by the nuanced task of analyzing short-form video transcripts. To summarize, our contributions are as follows: (i) the ViClaim dataset, a multilingual, multi-topic resource for claim detection in video transcripts; (ii) a custom annotation tool and comprehensive guidelines; (iii) baseline models to showcase the potential of ViClaim, and (iv) the open-sourcing of all resources, including the annotated transcripts with the raw annotations, video IDs, and experimental code.
@@ 3.6 Quality Control
  • FCW: Agreement is moderate on average ($0.4-0.59$), with Group 3 achieving substantial agreement ($0.64-0.76$).
  • FNC: Agreement is lower, with fair levels on average ($0.35-0.52$), though Group 3 achieved a maximum of $0.69$.
  • OPN: Agreement is moderate ($0.47-0.58$ on average), with Group 3 again demonstrating higher consistency ($0.58-0.66$).
  • None: This label consistently achieved the highest agreement, ranging from $0.57-0.85$, with averages close to substantial levels.
@@ 4.1 Experimental Setting
The models are trained to predict the labels of each sentence in a transcript. For this, let $\mathcal{C}=\{c_{i}\}_{i=1}^{1798}$ be the set of all clip transcripts. Each transcript is segmented into sentences that are labeled, thus $c_{i}=\{s_{j}^{(i)}\}_{j=1}^{n_{i}}$ where $n_{i}$ is the number of sentences of transcript $c_{i}$. Each sentence is a pair of text and label $s_{j}^{(i)}=(x_{j}^{(i)},y_{j}^{(i)})$, where $x_{j}^{(i)}=\{t_{k}\}_{k=1}^{m_{ij}}$ denotes the tokens ($m_{ij}$ number of tokens in sentence j of clip i), and $y_{j}^{(i)}$ is the label (cf. 3.6). The input to the classifiers is the concatenation of the full clip transcript for the context, with the sentence of the transcript to be classified, i.e., $\mathcal{D}=\{([x_{1}^{(i)};..;x_{n_{i}}^{(i)};x_{j}^{(i)}],y_{j}^{i})\}$, where $i$ denotes the clip number, $j^{th}$ the sentence within clip $i$ to be classified, and [;][;][ ; ] denotes the concatenation of sentence strings. During training, we use the soft labels as in Fornaciari et al. (2021) by computing the cross-entropy loss between the system output and the soft labels.
@@ 5 Discussion & Conclusion
In this study, we introduced ViClaim, a dataset of 1,798 annotated video transcripts spanning three languages (English, German, Spanish) and six topics, including socio-political and non-political content. The dataset employs a multi-label annotation approach, distinguishing between check-worthy and non-check-worthy facts and opinions. We fine-tuned state-of-the-art multilingual models, achieving strong performance in cross-validation (macro F1 up to 0.896). However, generalization to unseen topics remains challenging, particularly for non-political domains like League of Legends, highlighting the need for more advanced approaches to domain adaptation and contextual understanding in video-based misinformation detection. By releasing ViClaim and establishing strong baselines, we aim to drive progress in multimodal misinformation detection. This dataset addresses a critical research gap and serves as a foundation for developing tools to combat misinformation in the increasingly dominant medium of video.
@@
The annotation process encountered challenges due to the subjective nature of claim detection. Disagreement among annotators was not uncommon, as reflected in moderate Krippendorff’s $\alpha$ scores. Although MACE normalization was used to prioritize labels from more reliable annotators, this approach depends on the assumption that these annotators are consistently accurate across all contexts. Additionally, the multi-label annotation framework adds some complexity, especially for sentences with overlapping claims or mixed content, which should be considered when interpreting and applying the dataset to downstream tasks.
@@
We collected videos exclusively from YouTube and took care to comply with YouTube’s Terms of Service (ToS). Each video had a publicly visible URL at the time of selection and a median view count of 19097. We do not redistribute any copyrighted audio-visual content. Instead, we release only textual transcripts, treated as documented observations of publicly broadcast speech and their corresponding video IDs as reference. The original video materials remain on YouTube. Researchers wishing to access them must comply with YouTube’s ToS. Our dataset neither hosts nor reproduces the videos themselves, thereby avoiding copyright violations.
@@ Appendix A Hyperparameters
XLM-Robertal-Large (XLM). Conneau et al. (2020) pre-trained an 550M parameter encoder-transformer on 100 languages. For claim detection, we used a learning rate of $045e-$ 04, a batch size of 32 with a gradient accumulation of 8.
Falcon-7B (F7B). Almazrouei et al. (2023) pre-trained a 7B decoder-only model on 1.5T tokens of the RefinedWeb corpus Penedo et al. (2023). The main languages that Falcon performs well on are English, German, Spanish, and French, which cover our use case well. We applied QLorA Dettmers et al. (2024) with rank 16, alpha 32, a dropout of 0.05, and a learning rate of $045e-$ 04. For claim detection we used a batch size of 64 with a gradient accumulation of 8.
Mistral-7B (M7B). Jiang et al. (2023) pre-trained a 7B parameter decoder-only model. The details of the training data used are undisclosed. We applied QLorA Dettmers et al. (2024) with rank 16, alpha 8 a dropout of 0.05 and a learning rate of $045e-$ 04. For claim detection we used a batch size of 64 with a gradient accumulation of 8.
LLama3.2-3B (L3B). AI@Meta (2024) pre-trained an 8B parameter decoder-only model trained on an undisclosed set of 15T tokens of web data. We applied QLorA Dettmers et al. (2024) with rank 64, alpha 16, and dropout of 0.1, and a learning rate of $045e-$ 04. For claim detection we used a batch size of 64 with a gradient accumulation of 8.
@@ Appendix E Closed Source Models
In a preliminary step, we tested the performance of Grok2  [^12] https://x.ai/news/grok-2 on the classification task. For this, we prompted Grok2 with the task descriptions from our guidelines and added the examples from the paper to the prompt for in-context learning. The F1 scores are far below the fine-tuned models: F1 score for FCW: 0.665, F1 score for FNC: 0.549, and F1 score for OPN: 0.63. Thus, we abandoned closed-source models for this task and focused on fine-tuning classification models.

### v2

@@ 1 Introduction
In this work, we present the first step towards misinformation detection in videos by introducing ViClaim, a novel dataset where transcripts of YouTube Short videos are annotated using a custom annotation tool on a sentence level with the claim taxonomy introduced in Panchendrarajan and Zubiaga (2024). Each sentence is annotated to indicate whether it contains a claim, an opinion, or both, and if a claim is present, whether it is check-worthy. Since sentences may fall into multiple categories simultaneously, the task is framed as a multi-label classification problem. The dataset encompasses 1,798 videos in three languages (English, German, and Spanish) across six topics—five political and one entertainment—allowing for investigating domain transfer capabilities. Each sentence was annotated by four independent annotators, resulting in a total of 17,116 annotated sentences. Thus, ViClaim offers a rich database for future research on multimodal misinformation detection. As a first step to showcase the utility of ViClaim, we trained several baseline models on the transcripts (leaving multimodal models for future work). Our best-performing model achieved an F1 score of 89.9 for checkworthiness classification, 77.6 for detecting non-check-worthy claims, and 83.6 for opinion detection. These results demonstrate both the effectiveness of the dataset and the challenges posed by the nuanced task of analyzing short-form video transcripts. To summarize, our contributions are as follows: (i) the ViClaim dataset, a multilingual, multi-topic resource for claim detection in video transcripts; (ii) a custom annotation tool and comprehensive guidelines; (iii) baseline models to showcase the potential of ViClaim.
Release.
We release the corpus in form of the video IDs used, the time stamps of the annotated sentences, the corresponding labels, and the code pipeline to reconstruct the corpus, which can be found under the following GitHub repository  [^3] https://github.com/pgied/viclaim_stt. We release the experimental pipeline and the trained models, which can be found under the following GitHub repository  [^4] https://github.com/pgied/viclaim_training.
@@ 3.6 Quality Control
  • FCW: Agreement is moderate on average (0.40.590.4-0.59), with Group 3 achieving substantial agreement (0.640.760.64-0.76).
  • FNC: Agreement is lower, with fair levels on average (0.350.520.35-0.52), though Group 3 achieved a maximum of $0.69$.
  • OPN: Agreement is moderate (0.470.580.47-0.58 on average), with Group 3 again demonstrating higher consistency (0.580.660.58-0.66).
  • None: This label consistently achieved the highest agreement, ranging from 0.570.850.57-0.85, with averages close to substantial levels.
@@ 4.1 Experimental Setting
The models are trained to predict the labels of each sentence in a transcript. For this, let $\mathcal{C}=\{c_{i}\}_{i=1}^{1798}$ be the set of all clip transcripts. Each transcript is segmented into sentences that are labeled, thus $c_{i}=\{s_{j}^{(i)}\}_{j=1}^{n_{i}}$ where $n_{i}$ is the number of sentences of transcript $c_{i}$. Each sentence is a pair of text and label $s_{j}^{(i)}=(x_{j}^{(i)},y_{j}^{(i)})$, where $x_{j}^{(i)}=\{t_{k}\}_{k=1}^{m_{ij}}$ denotes the tokens ($m_{ij}$ number of tokens in sentence j of clip i), and $y_{j}^{(i)}$ is the label (cf. 3.6). The input to the classifiers is the concatenation of the full clip transcript for the context, with the sentence of the transcript to be classified, i.e., $\mathcal{D}=\{([x_{1}^{(i)};..;x_{n_{i}}^{(i)};x_{j}^{(i)}],y_{j}^{i})\}$, where $i$ denotes the clip number, $j^{th}$ the sentence within clip $i$ to be classified, and $[;]$ denotes the concatenation of sentence strings. During training, we use the soft labels as in Fornaciari et al. (2021) by computing the cross-entropy loss between the system output and the soft labels.
@@ 5 Discussion & Conclusion
In this study, we introduced ViClaim, a dataset of 1,798 annotated video transcripts spanning three languages (English, German, Spanish) and six topics, including socio-political and non-political content. The dataset employs a multi-label annotation approach, distinguishing between check-worthy and non-check-worthy facts and opinions. We fine-tuned state-of-the-art multilingual models, achieving strong performance in cross-validation (macro F1 up to 0.896). However, generalization to unseen topics remains challenging, particularly for non-political domains like League of Legends, highlighting the need for more advanced approaches to domain adaptation and contextual understanding in video-based misinformation detection. By releasing ViClaim and establishing strong baselines, we aim to drive progress in multi-modal misinformation detection. This dataset addresses a critical research gap and serves as a foundation for developing tools to combat misinformation in the increasingly dominant medium of video.
@@
The annotation process encountered challenges due to the subjective nature of claim detection. Disagreement among annotators was not uncommon, as reflected in moderate Krippendorff’s $\alpha$ scores. Although MACE normalization was used to prioritize labels from more reliable annotators, this approach depends on the assumption that these annotators are consistently accurate across all contexts. Additionally, the multi-label annotation framework adds some complexity, especially for sentences with overlapping claims or mixed content, which should be considered when interpreting and applying the dataset to downstream tasks. We note that our agreement scores show a medium agreement according to the standard interpretation. However, this does not imply that our annotations are of low quality, rather it highlights the ambiguity of our task. Thus, future research can work on modeling the uncertainty inherent in such a complex task.
@@
We collected videos exclusively from YouTube and took care to comply with YouTube’s Terms of Service (ToS). Each video had a publicly visible URL at the time of selection and a median view count of 19097. We do not redistribute any copyrighted audio-visual content. Instead, we release only the video IDs and the start-to-end timestamps for each annotated sentence. We also provide code that lets researchers recreate the transcripts from these references. The original video materials remain on YouTube. Researchers wishing to access them must comply with YouTube’s ToS. Our dataset neither hosts nor reproduces the videos themselves, thereby avoiding copyright violations.
@@ Appendix A Hyperparameters
XLM-Robertal-Large (XLM). Conneau et al. (2020) pre-trained an 550M parameter encoder-transformer on 100 languages. For claim detection, we used a learning rate of 5e045e-04, a batch size of 32 with a gradient accumulation of 8.
Falcon-7B (F7B). Almazrouei et al. (2023) pre-trained a 7B decoder-only model on 1.5T tokens of the RefinedWeb corpus Penedo et al. (2023). The main languages that Falcon performs well on are English, German, Spanish, and French, which cover our use case well. We applied QLorA Dettmers et al. (2024) with rank 16, alpha 32, a dropout of 0.05, and a learning rate of 5e045e-04. For claim detection we used a batch size of 64 with a gradient accumulation of 8.
Mistral-7B (M7B). Jiang et al. (2023) pre-trained a 7B parameter decoder-only model. The details of the training data used are undisclosed. We applied QLorA Dettmers et al. (2024) with rank 16, alpha 8 a dropout of 0.05 and a learning rate of 5e045e-04. For claim detection we used a batch size of 64 with a gradient accumulation of 8.
LLama3.2-3B (L3B). AI@Meta (2024) pre-trained an 8B parameter decoder-only model trained on an undisclosed set of 15T tokens of web data. We applied QLorA Dettmers et al. (2024) with rank 64, alpha 16, and dropout of 0.1, and a learning rate of 5e045e-04. For claim detection we used a batch size of 64 with a gradient accumulation of 8.
@@ Appendix E Closed Source Models
In a preliminary step, we evaluated three proprietary systems Grok2  [^14] https://x.ai/news/grok-2, o4-mini  [^15] https://platform.openai.com/docs/models/o4-mini and o3-mini  [^16] https://platform.openai.com/docs/models/o3-mini on our classification task. For this, we prompted each model with the task descriptions from our guidelines and added the examples from the paper to the prompt for in-context learning. Their F1 scores are far below the fine-tuned models listed in the table 7. The best performing models for each respecting label only achieved an F1 score for FCW: 0.7797, F1 score for FNC: 0.5559, and F1 score for OPN: 0.7789. Thus, we abandoned closed-source models for this task and focused on fine-tuning classification models.
F1
Model FCW FNC OPN
o3-mini 0.7797 0.5408 0.7789
4o-mini 0.6498 0.5559 0.7509
Grok 2 (closed source) 0.6650 0.5490 0.6300
Table 7: Scores for closed source models on the classification task.